OpenStudy (anonymous):

what is the reciprocal of 7 and 5/6

OpenStudy (senia01):

7 and 6/5 or 7/47

OpenStudy (albert0898):

A reciprocal is the number flipped.

OpenStudy (anonymous):

so it is D

OpenStudy (solomonzelman):

What is a reciprocal of: \(\large\color{black}{ \rm \displaystyle A+\frac{ b }{c} }\) ?? SOLUTION: At first, I will re-write the expression: \(\large\color{black}{ \rm \displaystyle A+\frac{ b }{c} =\frac{A\times c}{c}+\frac{b}{c} =\frac{A\times c+b}{c} }\) And then, yes, as you said reciprocal of a fraction is that fraction flipped. Recpriocal ───> \(\large\color{black}{ \rm \displaystyle\frac{c}{A\times c+b} }\)